Reeba has 17 socks in her drawer. 8 blues and 9 greens. She selects 5 socks at random. what is the probabilty that she gets:
a. 2 blues and 3 green
b. 3 blue and 2 green
c. at least one blue
d. no greens

Answer by stanbon(75887)

Reeba has 17 socks in her drawer. 8 blues and 9 greens. She selects 5 socks at random. what is the probabilty that she gets:
a. 2 blues and 3 green
# of ways to pick 2 blue = 8C2 = 8*7/1*2 = 28
# of ways to pick 3 green = 9C3 = 9*8*7/1*2*3 = 84
# of ways to pick 5 of any color = 17C5 =
-------
Therefore P(2 blue and 3 green) = [28*84]/17C5 = 0.38
----------------------------
b.P(3 blue and 2 green) = [8C3 * 9C2]/[17C5] = 0.32579...
---------------------
c.P(at least one blue) = 1 -P(no blue) = 1 - [9C5/17C5] = 0.9796...
---------------------
d.P(no greens) = 8C5/17C5
----------------
Cheers,
Stan H.